# On Neural Networks as Infinite Tree-Structured Probabilistic Graphical Models

### v3

@@ Appendix B HMC Trajectories
$\begin{split}\bm{\mu}_{n}(t+\frac{\Delta t}{2})&=\bm{\mu}_{n}(t)-\frac{\Delta t}{2}\nabla U(\bm{\rho})\bigg{|}_{\bm{\rho}=\bm{\rho}_{n}(t)}\\ \bm{\rho}_{n}(t+\Delta t)&=\bm{\rho}_{n}(t)+\Delta tM^{-1}\bm{\mu}_{n}(t+\frac{\Delta t}{2})\\ \bm{\mu}_{n}(t+\Delta t)&=\bm{\mu}_{n}(t+\frac{\Delta t}{2})-\frac{\Delta t}{2}\nabla U(\bm{\rho})\bigg{|}_{\bm{\rho}=\bm{\rho}_{n}(t+\Delta t)}.\end{split}$ (B.1)
@@
$\bm{h}^{(n+1)}\Big{|}(\bm{h}^{(n)}=\bm{\rho}_{n}(0))=\begin{cases}\bm{\rho}_{n}(L\Delta t)&\text{with probability $\alpha(\bm{\rho}_{n}(0),\bm{\rho}_{n}(L\Delta t))$}\\ \bm{\rho}_{n}(0)&\text{otherwise}\end{cases}$ (B.2)
@@
$\alpha(\bm{\rho}_{n}(0),\bm{\rho}_{n}(L\Delta t))=\min(1,\exp(H(\bm{\rho}_{n}(0),\bm{\mu}_{n}(0))-H(\bm{\rho}_{n}(L\Delta t),\bm{\mu}_{n}(L\Delta t)))).$ (B.3)
Appendix C Algorithm 1
Algorithm 1 CD-HMC Training for the Continuous Bernoulli Belief Network
1:Initialized $\bm{h}^{(0)}$, $\bm{W}^{(0)}=\{\bm{W}_{i}^{(0)},i=1,2,...,L\}$, $\bm{b}^{(0)}=\{\bm{b_{i}}^{(0)},i=1,2,...,L\}$

### v5

@@ Appendix E HMC Sampling Trajectories
$\begin{split}\bm{\mu}_{n}(t+\frac{\Delta t}{2})&=\bm{\mu}_{n}(t)-\frac{\Delta t}{2}\nabla U(\bm{\rho})\bigg{|}_{\bm{\rho}=\bm{\rho}_{n}(t)}\\ \bm{\rho}_{n}(t+\Delta t)&=\bm{\rho}_{n}(t)+\Delta tM^{-1}\bm{\mu}_{n}(t+\frac{\Delta t}{2})\\ \bm{\mu}_{n}(t+\Delta t)&=\bm{\mu}_{n}(t+\frac{\Delta t}{2})-\frac{\Delta t}{2}\nabla U(\bm{\rho})\bigg{|}_{\bm{\rho}=\bm{\rho}_{n}(t+\Delta t)}.\end{split}$ (E.1)
@@
$\bm{h}^{(n+1)}\Big{|}(\bm{h}^{(n)}=\bm{\rho}_{n}(0))=\begin{cases}\bm{\rho}_{n}(L\Delta t)&\text{with probability $\alpha(\bm{\rho}_{n}(0),\bm{\rho}_{n}(L\Delta t))$}\\ \bm{\rho}_{n}(0)&\text{otherwise}\end{cases}$ (E.2)
@@
$\alpha(\bm{\rho}_{n}(0),\bm{\rho}_{n}(L\Delta t))=\min\bigg{(}1,\exp\big{(}H(\bm{\rho}_{n}(0),\bm{\mu}_{n}(0)\big{)}-H(\bm{\rho}_{n}(L\Delta t),\bm{\mu}_{n}(L\Delta t))\big{)}\bigg{)}.$ (E.3)
Appendix F CD-k Algorithm
Algorithm 3 CD-$k$ Learning for the Deep Belief Network
1:Initialized $\bm{h}^{(0)}$, $\bm{W}^{(0)}=\{\bm{W}_{i}^{(0)}|i=1,2,...,K\}$, $\bm{b}^{(0)}=\{\bm{b_{i}}^{(0)}|i=1,2,...,K\}$
@@ Appendix F CD-k Algorithm
In our experiments, $k=1$, i.e. we do one sampling step before weight updating. We mainly use HMC sampling method for the sampling step, where the detailed trajectories are defined in Appendix E. The weight-updating step is defined in equations (2), (6) and (7).
We also run Gibbs method for comparison, where results are shown in Table 1. Both Gibbs and HMC are sampling methods to generate new states for all the hidden nodes and they both apply CD-1 or CD-k to learn the weight. While HMC samples real values in the range of $[0,1]$ under the normal distribution, Gibbs samples values in $\{0,1\}$ under Bernoulli distribution. For node $h_{ij}$, i.e., the $j$th node in $\bm{h}_{i}$ layer, the Bernoulli distribution used to sample its new state is determined by its Markov blanket, which includes all the nodes in $\bm{h}_{i-1},\bm{h}_{i},\text{and }\bm{h}_{i+1}$ in this case. It could be calculated by normalizing the joint probability distribution of the blanket when $h_{ij}=1$ versus $h_{ij}=0$, which is written as following:
$p(h_{ij}=1)=\frac{p(h_{ij}=1|\bm{h}_{i-1})\cdot p(\bm{h}_{i+1}|\bm{h}_{i}\backslash\{h_{ij}\},h_{ij}=1)}{p(h_{ij}=0|\bm{h}_{i-1})\cdot p(\bm{h}_{i+1}|\bm{h}_{i}\backslash\{h_{ij}\},h_{ij}=0)+p(h_{ij}=1|\bm{h}_{i-1})\cdot p(\bm{h}_{i+1}|\bm{h}_{i}\backslash\{h_{ij}\},h_{ij}=1)}$
We then sample a new state for $h_{ij}$ from $\text{Bern}(p(h_{ij}=1))$. In Gibbs sampling, the node is sampled one by one since a newly sampled node will affect the sampling of subsequent nodes in its Markov blanket. After sampling new values for all the hidden nodes, the weight is similarly updated using equations (2), (6) and (7). The Gibbs sampling does not depend on $L$.
Appendix G Experimental Setup
For all the experiments, the train-test split ratio is 80:20. For the training and finetuning, Adam optimizer is used with learning rate being $1\times 10^{-4}$. To get the predicted probabilities for fine-tuned network, 1000 output probabilities are sampled and averaged. In synthetic experiments, both BNs and MNs have the structure with the input dimension being 4, two latent layers with 4 nodes in each one, and one binary output.
Appendix H Running time
Running time of one BN experiment in the main text are shown in Table 3.
Table 3: Average running time of different methods on synthetic datasets. For DNN, it shows the time of the training for 100 and 1000 epochs. For Gibbs and HMC, it shows the time of the finetuning for 20 epochs.
Data (Weight) # Train Epochs Average Running Time (s)
DNN Gibbs HMC-10 HMC-100 HMC-1000
BN (0.3) 100 7.07 2769.81 666.33 650.10 758.37
1000 60.31 2741.30 634.20 626.79 728.57
BN (1) 100 7.22 2811.77 640.68 661.12 757.21
1000 65.25 2784.53 617.21 643.96 730.81
BN (3) 100 5.09 2616.04 615.96 615.05 636.93
1000 44.83 2602.65 596.23 594.62 608.02
BN (10) 100 6.87 2686.98 638.80 665.14 661.82
1000 48.98 2666.23 617.72 646.08 646.06
MN (0.3) 100 7.26 2703.71 645.57 635.42 729.60
1000 62.38 2669.61 625.38 618.07 701.31
MN (1) 100 4.94 2508.05 578.88 582.59 610.07
1000 44.39 2474.70 559.11 560.54 584.33
MN (3) 100 10.84 2610.14 591.43 579.26 605.13
1000 52.82 2578.81 570.23 558.98 585.22